OpenStudy (anonymous):

The volumes of two similar solids are 1,728 m^3 and 343 m^3. The surface area of the larger solid is 576 m^2. What is the surface area of the smaller solid? a. 196 m^2 b. 76m^2 c. 1,372m^2 d. 392 m^2

OpenStudy (anonymous):

(1728/343)^(1/3) = 12/7 A * (12/7)^2 = 576 So A = 196

OpenStudy (anonymous):

Thank you!(:

OpenStudy (anonymous):

Yup! Medal?

OpenStudy (anonymous):

How do you give medals? Sorry I'm new

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Can u help me with some more?